OpenStudy (anonymous):

Solve for x: 2 over 5x + 2 = 10

OpenStudy (anonymous):

2 over 5x as a decimal

OpenStudy (anonymous):

derp I mean fraction

OpenStudy (anonymous):

IS 2 over the whole equation or only 5x?

OpenStudy (anonymous):

only 5x

OpenStudy (anonymous):

so, 2/5x+2=10?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

okay, then x =1/20

OpenStudy (anonymous):

are you positive?

OpenStudy (anonymous):

ya, if the equation that I wrote is correct, then x=1/20

OpenStudy (anonymous):

shouldn't it be only 20? Because 1/20 isn't a choice

OpenStudy (anonymous):

Just plug in 1/20 into x and see if it equals 10

OpenStudy (anonymous):

10 isn't a choice either

OpenStudy (anonymous):

the answer is 1/20

OpenStudy (anonymous):



OpenStudy (anonymous):

Is that what your equation looks like?

OpenStudy (anonymous):

Yes yes it does

OpenStudy (anonymous):

ok, so then x is 1/20 or .05 as a fraction

OpenStudy (anonymous):

*decimal